OpenStudy (anonymous):

Solve: −negative 1 over 2 x + 3 = −x + 7 A) 1/2 B) 8 C) -8 D) 2

OpenStudy (geometry_hater):

Get x by itself

OpenStudy (anonymous):

yeah but what would the answer be?

OpenStudy (geometry_hater):

did you do the work yet? I cant just give you the answer......

OpenStudy (anonymous):

why not

OpenStudy (geometry_hater):

Because you'll never learn it You already know how you need to get x by itself.......Now what do you need to do first?

OpenStudy (anonymous):

okay is the answer 2?

OpenStudy (geometry_hater):

howd you get 2?

OpenStudy (anonymous):

1/2 + 3x = -x^2 + 7x x^2 -4x = -1/2 x^2 -4x + 4 = 4 - 1/2 (x-2)^2 = 14/4 x-2 = +/-sqrt(14)/2 x = [4 +/- sqrt14] /2

OpenStudy (geometry_hater):

There you go you showed your work and it is two btw good job

OpenStudy (anonymous):

thank you!